# Advances in Nonmonotone Proximal Gradient Methods merely with Local Lipschitz Assumptions in the Presense of Kurdyka–Łojasiewicz Property: A Study of Average and Max Line Search

### v1

@@ 4 NPG with Max Line Search and Convergence Analysis
Let Assumption 3.2 and Assumption 4.2,hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$, then the entire sequence $\{x^{k}\}$ converges to $\bar{x}$.
@@
By Proposition 4.3 (a), one knows that the whole sequence $\{q(x^{l(k)})\}$ is monotonically decreasing and convergent to $q(\bar{x})$ by assumption. This implies that $q(x^{l(k)})\geq q(\bar{x})$ for all $k\in\mathbb{N}$. If $q(x^{l(\bar{k})})=q(\bar{x})$ holds for some index $\bar{k}\in\mathbb{N}$, which, by monotonicity, implies that $q(x^{l(\bar{k}+1)})=q(\bar{x})$, then we claim that $x^{k+1}=x^{k}$ for all $k\geq\bar{k}+m_{\bar{k}}$ ($m_{\bar{k}}=m$ if $\bar{k}$ is sufficiently large) by [25, Lemma 3.2 (iv)]. Since, by assumption, the subsequence $\{x^{k}\}_{\mathcal{K}}$ converges to $\bar{x}$, this implies that $x^{k}=\bar{x}$ for all $k\in\mathbb{N}$ sufficiently large. In particular, we have the convergence of the entire (eventually constant) sequence $\{x^{k}\}$ to $\bar{x}$ in this situation.
@@
    $\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}},$
@@
$\|x^{l(k+1)}-x^{k+1}\|\leq\sum_{i=k+1-m}^{k+1}\|x^{i+1}-x^{i}\|\leq m\eta,\quad\forall k\geq k_{0},$ (4.19)
which means that $\|x^{l(k+1)}-x^{k+1}\|\to 0$. From (4.1), one has
$\gamma_{k}(x^{k}-x^{k+1})+\nabla f(x^{k+1})-\nabla f(x^{k})\in\hat{\partial}q(x^{k+1})\quad\forall k\in\mathbb{N}.$
Meanwhile, for all $k\in K$ where $K$ is defined in (4.4), we know that $x^{l(k+1)}\neq x^{k+1}$. We obtain that $\|x^{l(k+1)}-x^{k+1}\|\to 0$ by Theorem 4.7. Hence, by the definition of (Fréchet) subdifferential, we have
$q(x^{k+1})-q(x^{l(k+1)})-$ (4.20)
$\gamma_{l(k+1)-1}(x^{l(k+1)-1}-x^{l(k+1)})+\nabla f(x^{l(k+1)})-\nabla f(x^{l(k+1)-1}),x^{k+1}-x^{l(k+1)}\right>$
$\geq o(\|x^{k+1}-x^{l(k+1)}\|)$
Recall again that $x^{i}\in B_{\alpha}(\bar{x})$ for all $k_{0}\leq i\leq k+1$, then from (4.19), we have
@@
for all $k_{0}\leq i\leq k+1$. Therefore, for all $k_{0}\leq i\leq k+1$, we have $\gamma_{l(i+1)-1}\leq\bar{\gamma}_{\rho}$ from Lemma 4.4 that $\gamma_{l(i+1)-1}\leq\bar{\gamma}_{\rho}$, and $x^{l(i+1)-1},x^{l(i+1)}\in C_{\rho}$, then we have
$\|\nabla f(x^{l(i+1)})-\nabla f(x^{l(i+1)-1})\|\leq L_{\rho}\|x^{l(i+1)}-x^{l(i+1)-1}\|.$
Therefore, (4.20) implies that
$q(x^{i+1})-q(x^{l(i+1)})+(\bar{\gamma}_{\rho}+L_{\rho})\|x^{l(i+1)-1}-x^{l(i+1)}\|\|x^{i+1}-x^{l(i+1)}\|\geq o(\|x^{i+1}-x^{l(i+1)}\|)$
for all $k_{0}\leq i\leq k+1$ and $i\in K$. Employing the inequality $a^{2}+b^{2}\geq 2ab$ for all $a,b\in\mathbb{R}$ as well as $\|x^{i+1}-x^{l(i+1)}\|\to 0$ yields that
$q(x^{i+1})-q(x^{l(i+1)})+\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\left(\|x^{l(i+1)-1}-x^{l(i+1)}\|^{2}+\|x^{i+1}-x^{l(i+1)}\|^{2}\right)$
$\geq\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\|x^{i+1}-x^{l(i+1)}\|^{2}\quad\forall k_{0}\leq i\leq k+1\ \text{and}\ i\in K,$
which implies that
$\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\|x^{l(i+1)-1}-x^{l(i+1)}\|^{2}\geq q(x^{l(i+1)})-q(x^{i+1})$ (4.21)
for all $k_{0}\leq i\leq k+1$ and $i\in K$. By (4.21) and (4.4), we have
$\frac{\mu}{2}\|x^{i+1}-x^{i}\|^{2}<q(x^{l(i+1)})-q(x^{i+1})\leq\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\|x^{l(i+1)-1}-x^{l(i+1)}\|^{2}$
@@
$\|x^{i+1}-x^{i}\|\leq\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\|x^{l(i+1)-1}-x^{l(i+1)}\|\quad\forall k_{0}\leq i\leq k+1\ \text{and}\ i\in K.$
@@
$\sum_{k_{0}=i\in K}^{k+1}\|x^{i+1}-x^{i}\|\leq\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\sum_{k_{0}=i\in K}^{k+1}\|x^{l(i+1)-1}-x^{l(i+1)}\|$ (4.22)
$\leq\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\sum_{i=k_{0}}^{k+1}\|x^{l(i+1)-1}-x^{l(i+1)}\|$
$<\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\left(\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}\right).$
@@
$\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
$\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
@@
In particular, it follows from (a) that $x^{k}\in B_{\alpha}(\bar{x})$ for all $k\geq k_{0}$. Taking $k\to\infty$ in (4.14) therefore shows that $\{x^{k}\}$ is a Cauthy sequence and, thus, convergent. Since we already know that $\bar{x}$ is an accumulation point, it follows that the entire sequence $\{x^{k}\}$ is convergent to $\bar{x}$. ∎
Note that if $\overline{K}=\mathbb{N}$ or $m=0$, then we have $q(x^{k})>q(x^{k+1})$ for all $k\in\mathbb{N}$, i.e., $\{q(x^{k})\}$ is decreasing, then Algorithm 4.1 degenerates into [15, Algorithm 3.1], where the specific results about the convergence and convergence rate of the whole sequence were proposed. In the following, we illustrate the more general convergence rate of the whole sequence where $\overline{K}\neq\mathbb{N}$.
@@
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$. Then the entire sequence $\{x^{k}\}$ converges to $\bar{x}$, and if the corresponding desingularization function has the form $\chi(t)=ct^{\theta}$ for some $c>0$ and $\theta\in(0,1]$, then the following statements hold:
@@
    if $\theta=1$, then the sequences $\{q(x^{l(k)})\}$ and $\{x^{k}\}$ converge in a finite numer of steps to $q(\bar{x})$ and $\bar{x}$, respectively.
@@
    if $\theta\in[\frac{1}{2},1)$, then the sequence $\{q(x^{l(k)})\}$ converges Q-linearly to $q(\bar{x})$, and the sequence $\{x^{k}\}$ converges R-linearly to $\bar{x}$.
@@
Since $\{q(x^{l(k)})\}$ is decreasing, then statements (a), (b), and (c) regarding the sequence $\{q(x^{l(k)})\}$ follow from [14, Lemma 2.14]. It remains to verify the convergence rate with respect to the sequence $\{x^{k}\}$. We now consider the different cases of $\theta$.
  • $\theta=1$: for all $k\geq k_{0}$, (4.24) implies that $q(x^{l(k+1)})-q(x^{l(k)})\leq-\xi$, which yields that $\{q(x^{l(k)})\}$ converges to $q(\bar{x})$ in finite steps, then one has $\{x^{k}\}$ also converges to $\bar{x}$ in finite steps.
@@
    Note that $\{q(x^{l(k)})\}$ Q-linearly converges to $q(\bar{x})$, then $\{x^{k}\}$ R-linearly converges to $\bar{x}$ (by taking $s\to\infty$).

### v3

@@ 4 The NPG method with Max Line Search and Convergence Analysis
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$, then the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$.
@@
By Proposition 4.3 (a), one knows that the whole sequence $\{q(x^{l(k)})\}$ is monotonically decreasing and convergent to $q(\bar{x})$ by assumption. It implies that $q(x^{l(k)})\geq q(\bar{x})$ for all $k\in\mathbb{N}$. If $q(x^{l(\bar{k})})=q(\bar{x})$ holds for some index $\bar{k}\in\mathbb{N}$, which, by monotonicity, implies that $q(x^{l(\bar{k}+1)})=q(\bar{x})$, then we claim that $x^{k+1}=x^{k}$ for all $k\geq\bar{k}+m_{\bar{k}}$ ($m_{\bar{k}}=m$ if $\bar{k}$ is sufficiently large) by [26, Lemma 3.2 (iv)]. Since the subsequence $\{x^{k}\}_{k\in\mathcal{K}}$ is assumed to converge to $\bar{x}$, this implies that $x^{k}=\bar{x}$ for all $k\in\mathbb{N}$ sufficiently large. Then the entire (eventually constant) sequence $\{x^{k}\}_{k\in\mathbb{N}}$ is convergent to $\bar{x}$ in this situation.
@@
    $\leq\left(\frac{8\sqrt{2}m\left(c_{\hat{k}}+1\right)(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)$
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4\sqrt{2}m\left(c_{\hat{k}}+1\right)}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)$
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
@@
$\|x^{l(i+1)}-x^{i+1}\|\leq\sum_{j=i+1-m}^{i+1}\|x^{j+1}-x^{j}\|\leq m\eta,\quad\forall i\geq k_{0},$ (4.19)
which means that $\|x^{l(i+1)}-x^{i+1}\|\to 0$. By (4.1), for each $i\in\mathbb{N}$, one has
$\nabla f(x^{l(i+1)-1}),x^{l(i+1)}-x^{l(i+1)-1}\right>+\frac{\gamma_{l(i+1)-1}}{2}\|x^{l(i+1)}-x^{l(i+1)-1}\|^{2}+g(x^{l(i+1)})$
$\nabla f(x^{l(i+1)-1}),x^{i+1}-x^{l(i+1)-1}\right>+\frac{\gamma_{l(i+1)-1}}{2}\|x^{i+1}-x^{l(i+1)-1}\|^{2}+g(x^{i+1}),$
which implies that
$q(x^{l(i+1)})-q(x^{i+1})$ $\nabla f(x^{l(i+1)-1}),x^{i+1}-x^{l(i+1)}\right>+f(x^{l(i+1)})-f(x^{i+1})$ (4.20)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{\gamma_{l(i+1)-1}}{2}\left(\|x^{i+1}-x^{l(i+1)-1}\|^{2}-\|x^{l(i+1)}-x^{l(i+1)-1}\|^{2}\right).$
Recall again that $x^{i}\in B_{\alpha}(\bar{x})$ for all $k_{0}\leq i\leq k+1$, then $x^{i+1}\in C_{\rho}$ holds and from (4.19), we have
@@
for all $k_{0}\leq i\leq k+1$. Therefore, for all $k_{0}\leq i\leq k+1$, we have $\gamma_{l(i+1)-1}\leq\bar{\gamma}_{\rho}$ from Lemma 4.4, and $x^{i},x^{i+1},x^{l(i+1)-1},x^{l(i+1)}\in C_{\rho}$. Then the descent lemma implies
$f(x^{l(i+1)})-f(x^{i+1})\leq\frac{L_{\rho}}{2}\|x^{l(i+1)}-x^{i+1}\|^{2}+\left<\nabla f(x^{i+1}),x^{l(i+1)}-x^{i+1}\right>.$ (4.21)
Meanwhile, from the fact that $\frac{1}{2}\|\cdot-x^{l(i+1)-1}\|^{2}$ is convex, we have
$\frac{1}{2}\|x^{i+1}-x^{l(i+1)-1}\|^{2}-\frac{1}{2}\|x^{l(i+1)}-x^{l(i+1)-1}\|^{2}\leq\left<x^{i+1}-x^{l(i+1)-1},x^{i+1}-x^{l(i+1)}\right>$ (4.22)
Putting (4.20), (4.21), and (4.22) together yields that
$q(x^{l(i+1)})-q(x^{i+1})$ $\nabla f(x^{l(i+1)-1})-\nabla f(x^{i+1})+\gamma_{l(i+1)-1}\left(x^{i+1}-x^{l(i+1)-1}\right),x^{i+1}-x^{l(i+1)}\right>$ (4.23)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{L_{\rho}}{2}\|x^{l(i+1)}-x^{i+1}\|^{2}$
$\leq\frac{L_{\rho}}{2}\|x^{l(i+1)}-x^{i+1}\|^{2}+\left(\bar{\gamma}_{\rho}+L_{\rho}\right)\|x^{i+1}-x^{l(i+1)-1}\|\|x^{i+1}-x^{l(i+1)}\|$
$\leq\left(L_{\rho}+\frac{\bar{\gamma}_{\rho}}{2}\right)\|x^{l(i+1)}-x^{i+1}\|^{2}+\frac{\bar{\gamma}_{\rho}}{2}\|x^{l(i+1)}-x^{l(i+1)-1}\|^{2},$
therefore, by the inequality $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ for all $a\geq 0$ and $b\geq 0$, as well as Proposition 4.5, one has
$\sqrt{q(x^{l(i+1)})-q(x^{i+1})}$ $\leq\sqrt{L_{\rho}+\frac{\bar{\gamma}_{\rho}}{2}}\|x^{l(i+1)}-x^{i+1}\|+\sqrt{\frac{\bar{\gamma}_{\rho}}{2}}\|x^{l(i+1)}-x^{l(i+1)-1}\|$ (4.24)
$\leq\left(c_{\hat{k}}\sqrt{L_{\rho}+\frac{\bar{\gamma}_{\rho}}{2}}+\sqrt{\frac{\bar{\gamma}_{\rho}}{2}}\right)\|x^{l(i+1)}-x^{l(i+1)-1}\|$
for all $k_{0}\leq i\leq k+1$. Then, we have
$\|x^{i+1}-x^{i}\|<\sqrt{\frac{2}{\mu}}\sqrt{q(x^{l(i+1)})-q(x^{i+1})}\leq\left(c_{\hat{k}}\sqrt{\frac{2L_{\rho}+\bar{\gamma}_{\rho}}{\mu}}+\sqrt{\frac{\bar{\gamma}_{\rho}}{\mu}}\right)\|x^{l(i+1)-1}-x^{l(i+1)}\|$
@@
$\|x^{i+1}-x^{i}\|\leq\sqrt{2}\left(c_{\hat{k}}+1\right)\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\|x^{l(i+1)-1}-x^{l(i+1)}\|\quad\forall k_{0}\leq i\leq k+1\ \text{and}\ i\in K.$
@@
$\sum_{k_{0}=i\in K}^{k+1}\|x^{i+1}-x^{i}\|\leq\sqrt{2}\left(c_{\hat{k}}+1\right)\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\sum_{k_{0}=i\in K}^{k+1}\|x^{l(i+1)-1}-x^{l(i+1)}\|$ (4.25)
$\leq\sqrt{2}\left(c_{\hat{k}}+1\right)\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\sum_{i=k_{0}}^{k+1}\|x^{l(i+1)-1}-x^{l(i+1)}\|$
$\leq\frac{4\sqrt{2}m\left(c_{\hat{k}}+1\right)}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\left(\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}\right).$
@@
$\leq\left(\frac{8\sqrt{2}m\left(c_{\hat{k}}+1\right)(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4\sqrt{2}m\left(c_{\hat{k}}+1\right)}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
$\leq\left(\frac{8\sqrt{2}m\left(c_{\hat{k}}+1\right)(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4\sqrt{2}m\left(c_{\hat{k}}+1\right)}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
@@
In particular, it follows from (a) that $x^{k}\in B_{\alpha}(\bar{x})$ for all $k\geq k_{0}$. Taking $k\to\infty$ in (4.14) therefore shows that $\{x^{k}\}$ is a Cauthy sequence and, thus, convergent. Since we already know that $\bar{x}$ is an accumulation point, it follows that the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ is convergent to $\bar{x}$. ∎
Note that if $\overline{K}=\mathbb{N}$ or $m=0$, then we have $q(x^{k})>q(x^{k+1})$ for all $k\in\mathbb{N}$, i.e., $\{q(x^{k})\}_{k\in\mathbb{N}}$ is decreasing, then Algorithm 4.1 degenerates into [16, Algorithm 3.1], where the specific results about the convergence and convergence rate of the whole sequence were proposed. In the following, we illustrate the convergence rate of Algorithm 4.1 where $\overline{K}\neq\mathbb{N}$.
@@
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$. Then the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$, and if the corresponding desingularization function has the form $\chi(t)=ct^{\theta}$ for some $c>0$ and $\theta\in(0,1]$, then the following statements hold:
@@
    if $\theta=1$, then the sequences $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ and $\{x^{k}\}_{k\in\mathbb{N}}$ converge in a finite numer of steps to $q(\bar{x})$ and $\bar{x}$, respectively.
@@
    if $\theta\in[\frac{1}{2},1)$, then the sequence $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ converges Q-linearly to $q(\bar{x})$, and the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges R-linearly to $\bar{x}$.
@@
Since $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ is decreasing, then statements (a), (b), and (c) regarding the sequence $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ follow from [15, Lemma 2.14]. It remains to verify the convergence rate with respect to the sequence $\{x^{k}\}_{k\in\mathbb{N}}$. We now consider the different cases of $\theta$.
  • $\theta=1$: for all $k\geq k_{0}$, (4.27) implies that $q(x^{l(k+1)})-q(x^{l(k)})\leq-\xi$, which yields that $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ converges to $q(\bar{x})$ in finite steps, then one has $\{x^{k}\}_{k\in\mathbb{N}}$ also converges to $\bar{x}$ in finite steps.
@@
    Note that $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ Q-linearly converges to $q(\bar{x})$, then $\{x^{k}\}_{k\in\mathbb{N}}$ R-linearly converges to $\bar{x}$ (by taking $s\to\infty$).